write a system of two equation and use them to solve the following problem: two farmers were discussing how many chickens and cows they would purchase at market. the first farmer says to the second, " I have just purchased 30 new heads and 100 feet." how many chicken and how many cows did the farmer purchase.
Answer by Maths68(1474)


Let
Chickens = x
Cows = y
Thirty Heads
x+y=30
x=30-y...................(1)
Hundred Feet
2x+4y=100
2(x+2y)=100
2(x+2y)/2=100/2
x+2y=50...................(2)
Substitute the value of x from (1) to (2)
30-y+2y=50
30+y=50
y=50-30
y=20
put the value of y in (1)
x=30-y
x=30-20
x=10

Chickens = x = 10
Cows = y = 20